OpenStudy (anonymous):

solve the equation sec(θ)csc(θ)=2csc(θ) pls show steps

OpenStudy (anonymous):

divide = csc(x) sec(x) = csc(x) 1/sin(x) = 1/cos(x) sin(x) = cos(x) This is at 45degrees, and 225 degres

OpenStudy (anonymous):

Telling people to use wolfram alpha teaches nothing.

OpenStudy (anonymous):

they can see the whole stuff dear pk51. I understand it better with graphs and stuff. In any case, this is why many of us post answers

OpenStudy (anonymous):

trig gives me headaches no joke.

OpenStudy (anonymous):

when you divide by a trig function like that, just make sure you arent accidentally deleting one of the solutions. otherwise, nice nice. :)

myininaya (myininaya):

\[\sec(\theta)\csc(\theta)-2\csc(\theta)=0\] \[\csc(\theta)(\sec(\theta)-2)=0\]

OpenStudy (anonymous):

secx*cscx = 2cscx cosx*sinx = 1/2 *sinx 2sinx*cosx = sinx 2sinx*cosx - sinx = 0 sinx*(2cosx - 1) = 0 this gives sinx = 0 which is 0 or pi 2cosx - 1 = 0 cosx = 1/2 x = pi/3 or -pi/3

myininaya (myininaya):

joe i still do that sometimes and forget what i'm doing its the same reaon for if we have x^2=x if we divide both sides by x we are losing a solution x=1

OpenStudy (anonymous):

i got yelled at for making that mistake last semester lol. I'll never forget again >.<

myininaya (myininaya):

lol

OpenStudy (anonymous):

something about my solution is off....i guess i cant just flip all the functions? since its undefined if i do that?

OpenStudy (anonymous):

and lol >.< people cant take criticism =/ it was someone on here?

myininaya (myininaya):

yes

OpenStudy (anonymous):

wow this is like chinese to me what identities did you use because i dont recoginse it.

OpenStudy (anonymous):

do you know the ones like this? \[\sec(\theta) = \frac{1}{\cos(\theta)}, \csc(\theta) = \frac{1}{\sin(\theta)}\]

myininaya (myininaya):

here it was fine to divide by cscx since cscx is never zero :)

OpenStudy (anonymous):

\[\csc(\theta)(\sec(\theta)-2) = 0 \iff \csc(\theta) = 0, \sec(\theta)-2 = 0\]

OpenStudy (anonymous):

ah, cool cool, i see now.

OpenStudy (anonymous):

yes joe i do

OpenStudy (anonymous):

So at first i was trying to use those rules, but i ended up creating extra solutions by accident O.o the horror!

OpenStudy (anonymous):

but now i see, so csc can never be 0, so that means: \[\sec(\theta) - 2 = 0 \iff \sec(\theta) = 2 \iff \cos(\theta) = \frac{1}{2} \iff \theta = \frac{\pi}{6}, -\frac{\pi}{6}\]

OpenStudy (anonymous):

er...over 3, not 6 >.<

OpenStudy (anonymous):

\[\frac{π}{3},−\frac{π}{3}\]

OpenStudy (anonymous):

very nice drawing!

myininaya (myininaya):

:)

myininaya (myininaya):

i should had drawn her eating a duck

OpenStudy (anonymous):

i didnt know you could draw o.O

OpenStudy (anonymous):

myininaya that was excellent thx everyone too.

myininaya (myininaya):

i'm limited in my drawingness

OpenStudy (anonymous):

not as limited as i am lol >.< im only good at drawing ducks that look like birds.

myininaya (myininaya):

lol

myininaya (myininaya):

ok i gotta go take care of my cat for a few seconds

OpenStudy (anonymous):

alright, lates :P

OpenStudy (anonymous):

thx

myininaya (myininaya):

np